Disquotational Truth and Analyticity

Volker Halbach

The uniform reflection principle for the theory of uniform T-sentences

∀x(T A(ẋ) ↔ A(x))

is added to PA. The resulting system is justified on the basis of a disquotationalist theory of truth where the provability predicate is conceived as a special kind of analyticity; it is equivalent to the system ACA of arithmetical comprehension.

If the truth predicate is also allowed to occur in the instances of the T-sentences, yet not in the scope of negation, the system with the reflection schema for these stronger T-sentences assumes the strength of the Kripke-Feferman theory KF and thus to ramified analysis up to ϵ_{0}.

Address: Universität Konstanz, Fachgruppe Philosophie, Postfach 5560, 78434 Konstanz, Germany

E-mail: Volker.Halbach@uni-konstanz.de